# Rotatable Array-Aided Hybrid Beamforming for Integrated Sensing and Communication

### v1

@@ II-C Signal Model
Let $\mathbf{s}\in\mathbb{C}^{(K+1)\times 1}$ be the independent and identically distributed (i.i.d.) signal vector, with $\mathbb{E}(\mathbf{ss}^{\mathrm{H}})=\mathbf{I}_{\mathrm{K}+1}$, where $s_{1},\dots,s_{K}$ are for $K$ CUs, respectively, and $s_{K+1}$ is dedicated for sensing. Assume that the perfect channel state information (CSI) of communication channels is perfectly known at the BS via a proper channel estimation mechanism. The transimit beamforming matrice is
$\mathbf{W}=[\underbrace{\mathbf{w}_{1},\dots,\mathbf{w}_{K}}_{{\text{for communication and sensing}}},\underbrace{\mathbf{w}_{K+1}}_{{\text{dedicated for sensing}}}]\in\mathbb{C}^{N_{t}\times(K+1)},$ (13)
Then the received signal at the $k$-th user is then given by
@@
where $\mathbf{w}_{k}\in\mathbb{C}^{N_{t}\times\text{1}}$ denoted the digital transmit beamforming vector for the user $k\in\mathcal{K}$, $\mathbf{F}\in\mathbb{C}^{N_{t}\times N_{RF}}$ is the analog precoding matrix, and $n_{k}\sim\mathcal{C}\mathcal{N}(0,\sigma_{k}^{2})$ is zero-mean additive white Gaussian noise (AWGN) with noise power $\sigma_{k}^{2}$. With this sub-connected structure, the analog precoder matrix $\mathbf{F}$ is given by
@@
$\text{SINR}_{k}=\frac{|\mathbf{h}^{H}_{k}\mathbf{Fw}_{k}|}{\sum^{K+1}_{j=1,j\neq k}|\mathbf{h}^{H}_{k}\mathbf{Fw}_{j}|^{2}+\sigma_{k}^{2}},$ (17)
@@
where $n_{s}\sim\mathcal{C}\mathcal{N}(0,\sigma_{s}^{2})$ denotes the AWGN for radar link. $\zeta_{s}$ and $\zeta_{c}$ are complex coefficients including the radar cross section (RCS) of the target/ clutter, respectively. Additionally, $\mathbf{G}_{s}$ and $\mathbf{G}_{c}$ denote the array response vectors between the BS and the target/clutter, respectively.
We assume point-like target and clutters, since the sizes of the target and the clutters are suffciently small compared to the distances of the reflecting paths. Thus, the radar signal-to-clutter-plus-noise-ratio (SCNR) at the RX can be expressed as
@@
For point target detection in MIMO radar systems, to quantify how much information can be extracted with a given sensing signal, we can use the sensing mutual information (MI) per unit time [29][30]. It is defined as the MI between the received signal at the BS and the target response, conditioned on the transmitted signal. Specifically, the BS has knowledge of both the receive beamformer $\mathbf{u}$ and the transmitted ISAC signal x, the conditional sensing MI between the received echo and the target response (or the sensing rate) as [22]
@@ II-D Problem Formulation
where the weighting factors $\varpi_{s}$ and $\varpi_{c}$ control the priority of communication and sensing satisfy $\varpi_{s}+\varpi_{c}=1$, $\mathbf{\Lambda}=[\alpha^{\mathrm{t}},\beta^{\mathrm{t}},\gamma^{\mathrm{t}},\alpha^{\mathrm{r}},\beta^{\mathrm{r}},\gamma^{\mathrm{r}}]$ represent the sets of rotation angles for the transmit and receive planes. Constraint (21b) normalize the receive beamformer. Constraint (21c) denotes the transmit power should not exceed the maximum limits. Constraint (21d) corresponds to the sub-connected structure, i.e., the analog $\mathbf{F}$ belongs to a set of block matrices $\mathcal{A}_{p}$, where each block is an $M$ dimension vector with unit modulus elements. Due to the non-concave/non-convex objective function/constraints, problem (21) is difficult to solve. Specifically, in contrast to the existing FPA-based hybrid beamforming designs, the position change brought about by the rotatability of the transmitting and receiving antennas makes the variable coupling of the numerator and denominator in the achievable rates of users and targets more complex, thus posing further challenges. Moreover, Since the resource allocation needs to meet the different demands of the two functions, the trade-off between communication and sensing needs to be investigated.
It is worth mentioning that, for simplicity, we focus on a single target in this paper. Since our proposed scheme can treat the echoes from other targets as radar interference in the denominator of (19), the proposed algorithm can be similarly adapted to address the multi-target problem.
@@ III Proposed Algorithm
In this section, we propose an efficient algorithm for solving problem (21) by optimizing the transmit digitial beamformer, transmit analog beamformer, receive beamformer , the rotation angles of the transmitting and receiving plate. The problem is non-convex due to its objective function and coupling variables. To tackle this problem, we propose an alternating optimization (AO)-based algorithm where we update each variable with fixed values of the other variables obtained from the last iteration.
Given the mutual coupling between digital beamformer and analog beamformer in the objective and constraint (21c), it is generally difficult to optimize these two parameters simultaneously. Moreover, the non-convex SINR, SCNR and unit-modulus constraints make the problem even more challenging. A typical method is to first design the fully-digital beamformer and then approach the fully-digital one via hybrid beamformer design [9][31]. However, the obtained solution based on such a two-stage method cannot guarantee the original constraints to be satisfied. Therefore, we propose to directly optimize the digital beamformer and analog beamformer based on the principle of alternating optimization.
@@
For the variables $\mathbf{F}$, $\mathbf{W}$, $\mathbf{\Lambda}$, we employ the fractional programming (FP) approach [32]. We introduce auxiliary variables $\bm{\mu}=[\mu_{1},\dots,\mu_{K+1}]$, $\bm{\xi}^{c}=[\xi^{c}_{1},\dots,\xi^{c}_{K}]^{T}$ and $\bm{\xi}^{s}=[\xi^{s}_{1},\dots,\xi^{s}_{K+1}]^{T}$ to transform the (21a) into an equivalent convex form (22). Based on the AO framework, The problem (21) can be decomposed as five sub-problems. The details of the proposed algorithms are presented below.
@@ III-A Receive Beamforming Optimization
Given $\{\mathbf{F},\mathbf{W},\mathbf{\Lambda}\}$, the optimization of $\mathbf{u}$ only affect the SCNR (19). Therefore, maximizing the WSR, i.e., objective value (21a), is equivalent to maximizing SCNR (19). Hence, we formulate the subproblem for $\mathbf{u}$
@@
Based on the expressed in (23), it is found that the maximization of SCNR belongs to the problem of generalized Rayleigh quotient. By applying Proposition 1, the optimal closed-form solution can be obtained.
@@
we rewrite the numerator part of Formula (20) as
@@
Since $|\zeta_{s}|^{2}\mathbf{h}_{s}^{H}\tilde{\mathbf{A}}\mathbf{h}_{s}$ is non-negative and independent of the receive beamformer $\mathbf{u}$, based on the research results in Reference [33], we apply the generalized Rayleigh quotient to maximize problem (23) and arrive at the optimal solution $\mathbf{u}^{*}$, which is a minimal mean-square error (MMSE) filter. ∎
@@ III-B Transmit Digital Beamforming Optimization
Given $\{\mathbf{u},\mathbf{F},\mathbf{\Lambda},\bm{\mu},\bm{\xi}^{s},\bm{\xi}^{c}\}$, we optimize the digital beamformer $\mathbf{W}$. To decouple the product of $\mathbf{F}$ and $\mathbf{W}$ in the original power constraint (21c), we reformulate the power constraint in accordance with the special analog precoding structure as follows[]:
@@
Since $\tilde{\mathcal{G}}$ is a convex function with respect to $\mathbf{W}$, we can employ the Lagrange dual method to obtain the closed-form expression of $\mathbf{W}$. The Lagrangian function is defined as
@@ III-C Transmit Analog Beamforming Optimization
Note that the analog beamformer $\mathbf{F}$ is block diagonal, i.e., most of its element are zeros. To reduce the computational complexity, we extract the non-zero elements of $\mathbf{F}$ as
@@
where const($\bm{\mu}$) is a constant term when $\bm{\mu}$ is fixed, $\tilde{\mathbf{h}}_{k,j}^{H}=\mathbf{h}_{k}^{H}(\text{diag}(\mathbf{w}_{j})\otimes\mathbf{I}_{M})$, $\tilde{\bm{\beta}}_{k}^{H}=\sqrt{1+\mu_{k}}\,\xi_{k}^{c*}\,\tilde{\mathbf{h}}_{k,k}^{H}$, $\tilde{\bm{\beta}}_{s}^{H}=\sqrt{1+\mu_{K+1}}\zeta_{s}^{*}\mathbf{u}^{H}\mathbf{G}_{s}(\text{diag}(\mathbf{W}\bm{\xi}_{s})\otimes\mathbf{I}_{M})$ and $\tilde{\mathbf{G}}_{c,r}=\zeta_{c}\mathbf{G}_{c}(\text{diag}(\tilde{\mathbf{u}}_{r})\otimes\mathbf{I}_{M})$. Problem (41) is subject to the unit modulus constraint (41b). A conventional approach is to adopt Riemannian manifold optimization and solve it using the conjugate gradient descent method. However, this method suffers from low efficiency and unstable algorithm performance. To address the problem (41), we introduce the continuous variable $\phi$. Thus the subproblem (41a) is formulated as
@@
$\bm{\phi}^{*}=[\tilde{\bm{\Xi}}]^{-1}\tilde{\bm{\beta}}$ (43)
@@
$\text{arg}\{\mathbf{d}\}=\text{arg}\{\bm{\phi}\},$ (45)
Finally, by performing a simple matrix transformation on $\mathbf{d}$, the optimized variable $\mathbf{F}$ can be optained.
@@ III-D Auxiliary Variables Optimization
Similarly, we can derive the closed-form solution for the auxiliary variable $\bm{\mu}$ related to the lagrangian dual transform, which gives
@@ III-E Antenna Rotation Optimization
In this subsection, we optimize the antenna rotation angle $\bm{\Lambda}$ with any given other variables. $\mathbf{\Lambda}=[\alpha^{\mathrm{t}},\beta^{\mathrm{t}},\gamma^{\mathrm{t}},\alpha^{\mathrm{r}},\beta^{\mathrm{r}},\gamma^{\mathrm{r}}]$ is the set of rotation angles of the transmitting and receiving plates. To optimize $\mathbf{\Lambda}$, a feasible approach is to adopt an alternating optimization strategy, optimizing each element one by one until all elements in $\mathbf{\Lambda}$ are optimized. The transmit antenna rotation optimization and the receive antenna rotation optimization have similar mathematical formulations. For simplicity, we will focus on the optimization of the transmitting antenna rotation, and let $\mathbf{\Lambda}$ contain only one variable among $\alpha_{t}$, $\beta_{t}$ and $\gamma_{t}$, with this variable denoted as $\tilde{\Lambda}$. The subproblem of antenna rotation design is formulated as
@@
Problem (50) is difficult to solve directly because of its non-convexity. Here, we use $\tilde{\mathcal{G}}$ to replace $\tilde{\mathcal{G}}({\mathbf{u}},{\mathbf{W}},{\mathbf{F}},\bm{\mu},\bm{\xi}^{s},\bm{\xi}^{c})$ for simplicity. It is noted that $\tilde{\mathcal{G}}$ is differentiable across the entire feasible region, so we utilized the gradient ascent (GA) method to find a local stationary point efficiently. Derivative $\frac{\partial\tilde{\mathcal{G}}}{\partial\tilde{\Lambda}}$ can be rewritten as
@@
Define $\mathbf{D}_{k,n_{t}}=\mathbf{P}_{k}-\mathbf{t}_{n_{t}}$, $\bar{\mathbf{R}}_{v}$ denotes the derivative of the rotation matrix $\mathbf{R}$ with respect $v$, where $v\in\{\alpha,\beta,\gamma\}$. The partial derivatives w.r.t. $\tilde{\Lambda}$ are derived as (57)–(62).
@@
To address these shortcomings, we proposes an improved GA method. We first determine a relatively optimal initial point for GA through a preliminary search, and then employ an adaptive step size strategy for iterative searching until a feasible point is found or the convergence condition is met. Based on the above discussions, we propose an improved GA approach to solve the problem (50).
Initially, we set a discrete point set $\mathcal{K}$ with $K_{r}$ uniformly distributed points in the feasible region $\mathcal{N}$. Therefore, we can find an optimal initial $\tilde{\Lambda}$ that maximizes $\tilde{\mathcal{G}}$ as
@@
To ensure that in each iteration of the algorithm, the parameter is updated to a feasible point and remains within the feasible region, the update rule for $\tilde{\Lambda}^{(t+1)}$ is defined as:
@@
In this subsection, we focus on the case where $\mathbf{\Lambda}$ contains only one parameter. However, our method can be readily extended to scenarios involving multiple parameters. Specifically, it suffices to gradually optimize each variable in $\mathbf{\Lambda}$ using alternating iteration approach until the optimal $\mathbf{\Lambda}^{*}$ is obtained.

### v2

@@ II-C Signal Model
Let $\mathbf{s}\in\mathbb{C}^{(K+1)\times 1}$ be the independent and identically distributed (i.i.d.) signal vector, with $\mathbb{E}(\mathbf{ss}^{\mathrm{H}})=\mathbf{I}_{\mathrm{K}+1}$, where $s_{1},\dots,s_{K}$ are for $K$ CUs, respectively, and $s_{K+1}$ is dedicated for sensing. Assume that the perfect channel state information (CSI) of communication channels is perfectly known at the BS via a proper channel estimation mechanism. The transimit beamforming matrix is
$\mathbf{W}=[\underbrace{\mathbf{w}_{1},\dots,\mathbf{w}_{K}}_{{\text{for communication and sensing}}},\underbrace{\mathbf{w}_{K+1}}_{{\text{dedicated for sensing}}}]\in\mathbb{C}^{N_{t}\times(K+1)}.$ (13)
Then the received signal at the $k$-th user is given by
@@
where $\mathbf{w}_{k}\in\mathbb{C}^{N_{t}\times\text{1}}$ denotes the digital transmit beamforming vector for the user $k\in\mathcal{K}$, $\mathbf{F}\in\mathbb{C}^{N_{t}\times N_{RF}}$ is the analog precoding matrix, and $n_{k}\sim\mathcal{C}\mathcal{N}(0,\sigma_{k}^{2})$ is zero-mean additive white Gaussian noise (AWGN) with noise power $\sigma_{k}^{2}$. With this sub-connected structure, the analog precoder matrix $\mathbf{F}$ is given by
@@
$\text{SINR}_{k}=\frac{|\mathbf{h}^{H}_{k}\mathbf{Fw}_{k}|}{\sum^{K+1}_{j=1,j\neq k}|\mathbf{h}^{H}_{k}\mathbf{Fw}_{j}|^{2}+\sigma_{k}^{2}}.$ (17)
@@
where $n_{s}\sim\mathcal{C}\mathcal{N}(0,\sigma_{s}^{2})$ denotes the AWGN for radar link. $\zeta_{s}$ and $\zeta_{c}$ are complex coefficients including the RCS of the target/clutters, respectively. Additionally, $\mathbf{G}_{s}$ and $\mathbf{G}_{c}$ denote the array response vectors between the BS and the target/clutters, respectively.
We assume point-like target and clutter, since the sizes of the target and the clutter are sufficiently small compared to the distances of the reflecting paths. Thus, the radar signal-to-clutter-plus-noise-ratio (SCNR) at the RP can be expressed as
@@
For point target detection in MIMO radar systems, to quantify how much information can be extracted with a given sensing signal, we can use the sensing mutual information (MI) per unit time [30][31]. It is defined as the MI between the received signal at the BS and the target response, conditioned on the transmitted signal. Specifically, the BS has knowledge of both the receive beamformer $\mathbf{u}$ and the transmitted ISAC signal $\mathbf{x}$, the conditional sensing MI between the received echo and the target response (or the sensing rate) as [22]
@@ II-D Problem Formulation
where the weighting factors $\varpi_{s}$ and $\varpi_{c}$ control the priority of communication and sensing, and satisfy $\varpi_{s}+\varpi_{c}=1$. $\mathbf{\Lambda}=[\alpha^{\mathrm{t}},\beta^{\mathrm{t}},\gamma^{\mathrm{t}},\alpha^{\mathrm{r}},\beta^{\mathrm{r}},\gamma^{\mathrm{r}}]$ represents the sets of rotation angles for the transmit and receive panels. Constraint (21b) normalizes the receive beamformer. Constraint (21c) denotes the transmit power should not exceed the maximum power limit. Constraint (21d) corresponds to the sub-connected structure, i.e., the analog $\mathbf{F}$ belongs to a set of block matrices $\mathcal{A}_{p}$, where each block is an $M$ dimension vector with unit modulus elements. Due to the non-concave/non-convex objective function/constraints, problem (21) is difficult to solve. Specifically, in contrast to the existing FPA-based hybrid beamforming designs, the position changes caused by the rotatability of transmitting and receiving antennas complicate the variable coupling between the numerators and denominators in the achievable rates of users and targets. Moreover, since the resource allocation needs to meet the different demands of the two functions, the trade-off between communication and sensing needs to be investigated.
It is worth mentioning that, for simplicity, this paper focuses on a single target in this paper. Since the proposed algorithm can treat the echoes from other targets as radar interference in the denominator of (19), it can be similarly adapted to address the multi-target problem.
@@ III Proposed Algorithm
In this section, we propose an efficient algorithm for solving problem (21) by optimizing the transmit digital beamformer, the transmit analog beamformer, the receive beamformer, and the rotation angles of the transmitting and receiving panels. The problem is non-convex due to its objective function and coupling variables. To tackle this problem, an alternating optimization (AO) algorithm be proposed that updates each variable with fixed values of the other variables obtained from the last iteration.
Given the mutual coupling between the digital beamformer and the analog beamformer in the objective and constraint (21c), it is generally difficult to optimize these two parameters simultaneously. Moreover, the non-convex SINR, SCNR, and unit-modulus constraints make the problem even more challenging. A typical method is to first design the fully-digital beamformer and then approach the fully-digital one via hybrid beamformer design [9][32]. However, the obtained solution based on such a two-stage method cannot guarantee the original constraints to be satisfied. Therefore, we propose to directly optimize the digital beamformer and analog beamformer based on the principle of AO.
@@
For the variables $\mathbf{F}$, $\mathbf{W}$, $\mathbf{\Lambda}$, the fractional programming (FP) approach is employed[33]. We introduce auxiliary variables $\bm{\mu}=[\mu_{1},\dots,\mu_{K+1}]$, $\bm{\xi}^{c}=[\xi^{c}_{1},\dots,\xi^{c}_{K}]^{T}$ and $\bm{\xi}^{s}=[\xi^{s}_{1},\dots,\xi^{s}_{K+1}]^{T}$ to transform (21a) into an equivalent convex form (22). Based on the AO framework, the problem (21) can be decomposed as five sub-problems. The details of the proposed algorithms are presented below.
@@ III-A Receive Beamforming Optimization
Given $\{\mathbf{F},\mathbf{W},\mathbf{\Lambda}\}$, the optimization of $\mathbf{u}$ only affects the SCNR. Therefore, maximizing the WSR, i.e., objective value (21a), is equivalent to maximizing SCNR. Hence, the subproblem for $\mathbf{u}$ is formulated by
@@
Based on the expression in (23), the maximization of SCNR belongs to the problem of generalized Rayleigh quotient. By applying Proposition 1, the optimal closed-form solution can be obtained.
@@
Rewrite the numerator part of Formula (20) as
@@
Since $|\zeta_{s}|^{2}\mathbf{h}_{s}^{H}\tilde{\mathbf{A}}\mathbf{h}_{s}$ is non-negative and independent of the receive beamformer $\mathbf{u}$, based on the research results in Reference [34], the generalized Rayleigh quotient is applied to maximize problem (23) and arrive at the optimal solution $\mathbf{u}^{*}$, which is a minimal mean-square error (MMSE) filter. ∎
@@ III-B Transmit Digital Beamforming Optimization
Given $\{\mathbf{u},\mathbf{F},\mathbf{\Lambda},\bm{\mu},\bm{\xi}^{s},\bm{\xi}^{c}\}$, we optimize the digital beamformer $\mathbf{W}$. To decouple the product of $\mathbf{F}$ and $\mathbf{W}$ in the original power constraint (21c), we reformulate the power constraint in accordance with the sub-connected structure as follows:
@@
Since $\tilde{\mathcal{G}}$ is a convex function with respect to $\mathbf{W}$, the Lagrange dual method can be employed to obtain the closed-form expression of $\mathbf{W}$. The Lagrangian function is defined as
@@ III-C Transmit Analog Beamforming Optimization
Note that the analog beamformer $\mathbf{F}$ is block diagonal, i.e., most of its elements are zeros. To reduce the computational complexity, we extract the non-zero elements of $\mathbf{F}$ as
@@
where const($\bm{\mu}$) is a constant term when $\bm{\mu}$ is fixed, $\tilde{\mathbf{h}}_{k,j}^{H}=\mathbf{h}_{k}^{H}(\text{diag}(\mathbf{w}_{j})\otimes\mathbf{I}_{M})$, $\tilde{\bm{\beta}}_{k}^{H}=\sqrt{1+\mu_{k}}\,\xi_{k}^{c*}\,\tilde{\mathbf{h}}_{k,k}^{H}$, $\tilde{\bm{\beta}}_{s}^{H}=\sqrt{1+\mu_{K+1}}\zeta_{s}^{*}\mathbf{u}^{H}\mathbf{G}_{s}(\text{diag}(\mathbf{W}\bm{\xi}_{s})\otimes\mathbf{I}_{M})$ and $\tilde{\mathbf{G}}_{c,r}=\zeta_{c}\mathbf{G}_{c}(\text{diag}(\tilde{\mathbf{u}}_{r})\otimes\mathbf{I}_{M})$. Problem (41) is subject to the unit modulus constraint (41b). A conventional approach is to adopt Riemannian manifold optimization and solve it using the conjugate gradient descent method. However, this method suffers from low efficiency and unstable algorithm performance. To address the problem (41), the continuous variable $\phi$ is introduced. Thus the subproblem (41a) is formulated as
@@
$\bm{\phi}^{*}=[\tilde{\bm{\Xi}}]^{-1}\tilde{\bm{\beta}}$, (43)
@@
$\text{arg}\{\mathbf{d}\}=\text{arg}\{\bm{\phi}\}.$ (45)
Finally, by performing a simple matrix transformation on $\mathbf{d}$, the optimized variable $\mathbf{F}$ can be obtained.
@@ III-D Auxiliary Variables Optimization
Similarly, the closed-form solution for the auxiliary variable $\bm{\mu}$ related to the lagrangian dual transform, which is given by
@@ III-E Antenna Rotation Optimization
In this subsection, we optimize the antenna rotation angle $\bm{\Lambda}$ with any given other variables. $\mathbf{\Lambda}=[\alpha^{\mathrm{t}},\beta^{\mathrm{t}},\gamma^{\mathrm{t}},\alpha^{\mathrm{r}},\beta^{\mathrm{r}},\gamma^{\mathrm{r}}]$ is the set of rotation angles of the transmitting and receiving panels. To optimize $\mathbf{\Lambda}$, a feasible approach is to adopt an alternating optimization strategy, optimizing each element one by one until all elements in $\mathbf{\Lambda}$ are optimized. The transmit antenna rotation optimization and the receive antenna rotation optimization have similar mathematical formulations. For simplicity, this paper focuses on the optimization of the transmitting antenna rotation, and let $\mathbf{\Lambda}$ contain only one variable among $\alpha_{t}$, $\beta_{t}$ and $\gamma_{t}$, with this variable denoted as $\tilde{\Lambda}$. The subproblem of antenna rotation design is formulated as
@@
Problem (50) is difficult to solve directly because of its non-convexity. Here, $\tilde{\mathcal{G}}({\mathbf{u}},{\mathbf{W}},{\mathbf{F}},\bm{\mu},\bm{\xi}^{s},\bm{\xi}^{c})$ is replaced by $\tilde{\mathcal{G}}$ for simplicity. It is noted that $\tilde{\mathcal{G}}$ is differentiable across the entire feasible region, so the gradient ascent (GA) method be utilized to find a local stationary point efficiently. Derivative $\frac{\partial\tilde{\mathcal{G}}}{\partial\tilde{\Lambda}}$ can be rewritten as
@@
Define $\mathbf{D}_{k,n_{t}}=\mathbf{P}_{k}-\mathbf{t}_{n_{t}}$. $\bar{\mathbf{R}}_{v}$ denotes the derivative of the rotation matrix $\mathbf{R}$ with respect $v$, where $v\in\{\alpha,\beta,\gamma\}$. The partial derivatives w.r.t. $\tilde{\Lambda}$ are derived as (57)–(62).
@@
To address this issue, we first determine a relatively optimal initial point for GA through a preliminary search, and then employ an adaptive step size strategy for iterative searching until a feasible point is found or the convergence condition is met. Based on the above discussions, an improved GA approach is proposed to solve the problem (50).
Initially, we set a discrete point set $\mathcal{K}$ with $K_{r}$ uniformly distributed points in the feasible region $\mathcal{N}$. Therefore, an optimal initial $\tilde{\Lambda}$ can be find that maximizes $\tilde{\mathcal{G}}$ as
@@
To ensure that in each iteration of the algorithm, the parameter is updated to a feasible point and remains within the feasible region. The update rule for $\tilde{\Lambda}^{(t+1)}$ is defined as:
@@
In this subsection, we focus on the case where $\mathbf{\Lambda}$ contains only one parameter. However, our method can be readily extended to scenarios involving multiple parameters. Specifically, it suffices to gradually optimize each variable in $\mathbf{\Lambda}$ using an alternating iteration approach until the optimal $\mathbf{\Lambda}^{*}$ is obtained.